Find the accumulated amount after 9 yr if $3900 is invested at 3%/year compounded continuously.
Answer by rfer(16322)

A=Pe^rt
A=3900*(2.7183)^0.03*9
A=3900*(2.7183)^1.06814
A=3900*$11,348.93